# Prompt Stellar and Binary Black Hole Mergers from Chemically Homogeneous Evolution in Triples

### v1

@@ 4.1.2 Binary Black Holes
Our default CHE binary evolves effectively undisturbed if the tertiary is $m_{3}<55\ M_{\odot}$ and the outer orbital period is $P_{\rm out}\gtrsim 100$ d (Fig. 7). In this regime, the inner CHE binary does not merge, but leads to the formation of a BBH. Once a BBH is formed, tidal forces are no longer present and the GR precession becomes the only effect suppressing the ZLK mechanism. This may result in triples where the ZLK effect can significantly perturb the eccentricity of the BBH. In our default CHE binary, in certain triple configurations, the time-to-coalescence via gravitational-wave emission can decrease by several orders of magnitude, leading to prompt mergers only $\sim 100$ yr after BBH formation (Fig. 7). This decrease is even more prominent in the alternative CHE binary at higher $Z=0.0042$. At that metallicity, the BBH would not merge within the age of the Universe unless dynamically assisted by a tertiary (§3.3), and its merger timescale can be of the order of the ZLK timescale (Fig. 8).
The effects of a tertiary companion around a BBH might be varied and are uncertain. Dorozsmai et al. (2024) demonstrated that triples with CHE binaries can experience tertiary mass transfer onto a BBH. Subsequently, Kummer et al. (2025) examined these configurations in greater detail, discovering that gas accretion onto the inner BBH can facilitate drag-assisted mergers, potentially resulting in an electromagnetic counterpart. In addition, here we find that some BBH mergers can promptly occur purely through dynamical assistance via the ZLK mechanism. When a BBH merges, it losses some mass-energy through gravitational waves, and depending on the spin configuration, it may receive a recoil kick. BBHs formed from CHE evolution are believed to have notably high spins, unlike the typically non-rotating black holes from other evolutionary channels in isolated binary evolution (Marchant et al., 2024). For isolated CHE binaries, these spins are thought to be aligned; however, it is not clear how black hole formation and precession might affect the direction of these spins (e.g., Tauris, 2022). Depending on the alignment of these spins, the recoil kick for the black hole remnant could range from $\sim 10-1000$ km/s. Such a recoil kick could propel the black hole through surrounding material lost via winds or during black hole formation (e.g., Burrows et al., 2023) during the late stages of evolution, as well as prompt interactions with the tertiary if it is a stellar companion. Therefore, we highlight these ZLK-assisted BBH mergers as potential candidates for joint gravitational-wave and electromagnetic detection.
@@ 4.2.1 Stellar Evolution
Our analysis examined two stellar models with the same mass but different metallicities. The evolution of massive stars is influenced by various uncertainties, including convection, mixing, mass loss, and magnetic fields. Here we emphasize two main caveats: stellar expansion and collapse.
@@
In this paper, we focused on the CHE scenario, where rotational mixing allows the star to remain compact throughout the main sequence and leads to contraction afterward. Vigna-Gómez et al. (2021) note that any compact star can play a significant role in the evolution of tight triples, including stars in metal-free (Population III) environments (e.g., Marigo et al., 2001) or within specific mass–metallicity regimes (e.g., Shenar et al., 2020). We adhere this idea and propose that prompt (stellar and BBH) mergers are likely to occur in systems where the component stars of the inner binary do not significantly expand after the main sequence, regardless of the physical mechanism behind this.
With respect to stellar collapse, we assume that our $\approx 43.8\ M_{\odot}$ carbon-oxygen core will form a black hole. We further assume that this black hole is created through direct collapse, without any baryonic mass ejection or associated explosion. Energy is lost exclusively via neutrinos, resulting in a mass decrement of $\sim 0.1\ M_{\odot}$ and no natal kick, consistent with some observed systems (e.g., Vigna-Gómez et al., 2024, and references therein). However, it is important to acknowledge the possibility of mass ejection during black hole formation, as well as a natal kick. Mass ejection could contribute to the energy reservoir relevant for an electromagnetic counterpart to the BBH merger (e.g., de Mink & King, 2017; Burrows et al., 2023). A natal kick could affect the orbits and their dynamics. These effects are not considered in this study.
@@ 4.2.2 Binary Evolution
In our analysis, the role of tides in the inner binary is crucial (§2.3). Tides influence the evolution of binary and triple systems through various effects, such as rotation, tidal distortion, and tidal dissipation (§2.3.4).
@@
Besides rotation, the presence of a close companion can further tidally distort the star, enhancing mixing efficiency compared to models that consider only rotation (Hastings et al., 2020). Additionally, in a tidally deformed short-period binary a non-dissipative tidal bulge may form. This bulge can lead to apsidal precession, which suppresses the ZLK mechanism (§2.2.2). Our default CHE binary is not only a rotating, short-period binary but also a contact system. In contact binaries, tidal distortion can cause a radius discrepancy of up to 5% (Fabry et al., 2022); however, this discrepancy does not significantly impact our calculations.
In our framework for estimating the maximum eccentricity (§2.2.2), we assume that tides influence the system through the potential created by the non-dissipative tidal bulge, with no energy dissipation occurring in the system. However, in binaries with orbital periods less than approximately 10 days, tides serve as an efficient mechanism for energy dissipation. In this paper, we focused on the dynamical tide, which is suitable for stars with a radiative envelope, as the primary tidal dissipation mechanism (§2.3); we estimated $\tau_{\rm{circ}}\approx 1500\ \rm yr$ and $\tau_{\rm{dyn}}\approx 23\ \rm yr$. Within the formalism for the dynamical tide, there are major uncertainties in the value of the second-order tidal coefficient $E_{2}$ (e.g., Siess et al., 2013; Kushnir et al., 2017), which is inversely proportional to $\tau_{\rm{circ}}$ (Eq. (12)). Different models for of $E_{2}$ can lead to shorter ($\tau_{\rm{circ}}\approx 185$ yr and $\tau_{\rm{dyn}}\approx 3$ yr, Qin et al. 2018) or longer ($\tau_{\rm{circ}}\approx 2400$ yr and $\tau_{\rm{dyn}}\approx 37$ yr, Yoon et al. 2010) estimates for the tidal timescale at the ZAMS and throughout the stellar lifetime. These different parametrisations can either constrain or expand the parameter space where the ZLK mechanism influences the eccentricity of the inner binary.
@@ 4.2.3 Triple Evolution
In the study of multiple star systems, it is often assumed that the component stars are coeval. This assumption was applied in our analysis. However, simulations indicate that some triples form after the inner binary has assembled, either as part of the same starburst (e.g., Bonnell & Bate, 1994) or through a dynamical capture. In our tight triple systems, the ZLK timescale ($\tau_{\rm ZLK}\lesssim 10^{4}$ yr) is very short with respect to the stellar lifetime ($\sim 10^{6}$ yr). Consequently, for a coeval triple system, a merger near the ZAMS would effectively correspond to a merger in the pre-main-sequence phase. Alternatively, the merger might occur at a much later stage, after the triple system has formed and achieved a dynamical steady-state equilibrium. Only then would it evolve in a manner more akin to the scenario considered in our formalism. In such cases, the merger would not occur around the ZAMS but shortly after the triple assembled.
Once the triple system formed, there are additional considerations to keep in mind. In our analysis, we have assumed that the mass ratio is close to one and that the outer orbit is circular. This assumption for the mass ratio is common in the literature on CHE binaries (Marchant et al., 2016; Mandel & de Mink, 2016). However, massive contact binaries are often observed with mass ratios that differ from unity (see the sample presented in Menon et al., 2021). While Fabry et al. (2023) determined that energy transfer in massive contact binaries alters the evolution of the mass ratio and can prolong the lifespan of a contact system with unequal masses, Fabry et al. (2024) showed that this effect is small at a population level.
The assumption about a circular outer orbit might not be representative for hierarchical triple systems (e.g., Hwang et al., 2022). The two tight high-mass triples, TIC 470710327 (Eisner et al., 2022) and TIC 290061484 (Kostov et al., 2024), have outer eccentricities of $e_{\rm out}=0.3$ and $e_{\rm out}=0.2$, respectively. Also, the tertiary around a massive binary may have an eccentric orbit due to the apsidal precession resonance (e.g., Liu & Lai, 2024). For systems with non-equal mass binaries and non-zero outer eccentricity, the octupole effect may become important (Naoz, 2016). This tends to widen the inclination window for large eccentricity excitation. However, the analytic expression for $e_{\rm max}$ given by Eq. (6) remains valid. Thus, because of the effect of short-range forces due to tides, the maximum eccentricity cannot exceed $e_{\rm max}$ even when the octupole potential is significant (Liu et al., 2015).
@@ 4.3 Constraining Tidal Theory
There are significant uncertainties in our understanding of tidal theory, particularly in massive, short-period binary systems. These uncertainties primarily stem from the standard model of the dynamical tide (Zahn, 1975, 2008), which describes how stellar oscillation modes are excited by a periodic tidal potential. A key parameter in this model, the $E_{2}$ coefficient (§2.3), is highly sensitive to the definition of the core-envelope boundary (e.g., Siess et al., 2013; Kushnir et al., 2017), making it a crucial but uncertain factor in tidal evolution.
@@
The interplay between the inner and outer binaries remains highly uncertain. Short-range forces, including rotation, tides, and GR effects, have been shown to suppress the ZLK mechanism (Liu et al., 2015). In our default CHE binary model, tides are the dominant short-range force responsible for this suppression. However, the structure and evolution of the tidal bulge—which determines the maximum eccentricity of the inner binary and whether it will merge as a stellar or BBH system—are still not well understood.
To refine tidal theory, we propose using future observations of high-mass tight triple systems. The mere detection of such systems could provide empirical constraints on tidal effects. If the orbital configuration suggests that the ZLK mechanism should be active, we can compare cases with and without tidal suppression (ZLK+SA+GR+Tides vs ZLK+SA+GR) to determine whether tides are strong enough to prevent eccentricity growth. If the inner binary remains circular over the ZLK timescale (§2.3), this would place an upper limit on the tidal dissipation timescale. Additionally, such systems could allow us to solve for $E_{2}$ in specific binary configurations, further improving tidal models. Conversely, if the inner binary exhibits eccentricity evolution, we can compare the observed evolution to our ZLK model to assess the accuracy of our tidal prescriptions, either by determining the value of $\epsilon_{\rm Tide}$ or by quantifying deviations from the expected ZLK+SA+GR solution.
In summary, observations of tight triple systems offer a direct way to test and refine tidal theory. By analyzing their orbital configurations and evolutionary pathways, we can place new constraints on tidal dissipation, improve our understanding of how tides suppress eccentricity growth, and ultimately enhance our models of stellar and compact object mergers.
@@ 4.4 Population Forecasting for Tight Triples
Finally, tight triples in which the inner binary forms a BBH present very interesting prospects. These can result in triple systems with a tertiary in an outer orbital period of 100 or 1000 d, detectable with facilities such as, e.g., Gaia (Gaia Collaboration et al., 2016). Triple configurations that lead to a prompt BBH merger could potentially be detected by the LIGO-Virgo-Kagra detector network, and result in a binary configuration of an unusually massive black hole orbited by a nearby stellar companion.
@@ 5 Summary and Conclusions
We investigated the role of the von Zeipel-Lidov-Kozai (ZLK) mechanism in high-mass triple systems. We focused on ‘tight’ triples, which are composed of a short-period, high-mass stellar binary with a nearby tertiary companion. We explored an illustrative case involving a circular binary with component masses $m_{1}=m_{2}=55\ M_{\odot}$, initial orbital period $P_{\rm in}=1.1$ d at ‘low’ ($Z=0.00042$) and ‘moderately-low’ ($Z=0.0042$) metallicities (§2.1 and Appendix A, respectively). This binary was simulated using the detailed stellar evolution software MESA.
@@
The component stars in this binary undergo chemically homogeneous evolution (CHE), resulting in well-mixed stars that remain compact throughout the main sequence and contract afterward (Fig. 1 and 9). Based on our numerical model, we examined the impact of a tertiary with a mass range between $1\leq m_{3}/M_{\odot}\leq 100$ and an outer orbital period range between $1\leq P_{\rm out}/\rm{d}\leq 1000$, assuming the initial mutual inclination is uniformly distributed in the range $-1\leq\cos(i_{0})\leq 1$.
@@
We utilized the analytic results from Liu et al. (2015) and Mangipudi et al. (2022) to determine the maximum eccentricity that the inner binary can achieve in different triple configurations (§2.2 and Fig. 5). We evaluated whether the ZLK mechanism could cause the component stars of the binary to fill their outer Lagrangian points. If this occurs, we consider the inner binary to result in a stellar merger (§3.2). We find three key evolutionary stages for mergers: near the zero-age main sequence, shortly after the end of the main sequence, of after the formation of a binary black hole (BBH, Fig. 7 and 8).
We finalize with our concluding remarks.
@@
  • Main-sequence Mergers. During the main sequence, strong tidal bulges in close binaries significantly damp the maximum eccentricity produced by the ZLK mechanism (Fig. 3 and 5). Nonetheless, mergers can occur even at moderately-low eccentricities ($e\gtrsim 0.3$), with an occurrence fraction of $f_{\rm{merger}}\approx 0.77$ for optimal inclinations. These mergers are possible in systems with outer periods of up to $\approx 70$ d at low metallicity (Fig. 6) and up to $\approx 100$ d at moderately-low metallicity (Fig. 10). Given the ZLK timescales (Fig. 4), these hydrogen-rich mergers must happen near the zero-age main sequence for coeval triples or shortly after the formation of the triple for non-coeval systems.
@@
  • Post-main-sequence Mergers. After the main sequence, CHE stars rapidly contract and the effect of tides weakens. At this stage, mergers can occur at high eccentricities ($e\gtrsim 0.8$), and occur less frequently at a fraction of $f_{\rm merger}\approx 0.57$ for optimal inclinations. These mergers are helium-rich and can occur in systems with outer periods of up to $\approx 100$ d at low metallicity (Fig. 6) and up to $\approx 270$ d at moderately-low metallicity (Fig. 10).
@@
  • Binary Black Hole Mergers. For triples in which the inner binary does not undergo a stellar merger, a BBH may form. In BBHs, tides are no longer present, and the ZLK mechanism is largely suppressed by general relativity effects. The merger of BBHs can be facilitated by a tertiary companion with an orbital period of up to a few thousand days, and in favourable orbital configurations the time-to-coalescence may be as short as the ZLK timescale (Fig. 7 and 8). This results in prompt mergers shortly after the formation of a BBH within a tight triple system, which contrasts with the generally slower coalescence expected for BBHs formed in isolation. These BBH mergers occur near recently ejected gas —from stellar winds or following core collapse— and possibly close to another star. This makes them promising candidates for electromagnetic counterparts and probes of environmental effects in gravitational-wave sources.
@@
  • Orbital Configurations Leading to Mergers. Contrary to the test particle limit, maximum eccentricity is not always achieved at a $90$ deg mutual inclination. Instead, it depends on the ratio of the inner to outer orbital angular momenta ($\eta:=L_{\rm in}/L_{\rm out}$, Fig. 2). Moreover, when the tertiary is less massive than the inner component stars of the binary, the maximum eccentricity is reached at largely retrograde configurations (Fig. 6 and 10). This behaviour is independent of the nature of the component stars of the inner binary; it should also occur in stars that do not evolve via CHE.
@@
  • Outlook. The dynamics of tight, high-mass stellar triples are complex. When the tertiary star is more massive than either component stars of the inner binary, the evolution of the system approximates the test particle limit. However, in this regime, stellar evolution —particularly mass transfer from the tertiary— can play a crucial role. Conversely, if the tertiary is less massive than either inner companions, evolution is dominated by the inner binary, and even significantly inclined systems may experience reduced dynamical effects due to being far from the test particle limit. On a population level, an important question is which tight triples reduce to binaries and which remain intact as triples, the latter being candidate progenitors of merging BBHs.

### v2

@@ 4.1.2 Binary black holes
Our default CHE binary evolves effectively undisturbed if the tertiary is $m_{3}<55\ M_{\odot}$ and the outer orbital period is $P_{\rm out}\gtrsim 100$ d (Fig. 7). In this regime, the inner CHE binary does not merge, but leads to the formation of a BBH. Once a BBH is formed, tidal forces are no longer present and the GR precession becomes the only effect suppressing the ZLK mechanism. This may result in triples where the ZLK effect can significantly perturb the eccentricity of the BBH. In our default CHE binary, in certain triple configurations, the time-to-coalescence via gravitational-wave emission can decrease by several orders of magnitude, leading to prompt mergers only $\sim 100$ yr after BBH formation (Fig. 7). This decrease is even more prominent in the alternative CHE binary at higher $Z=0.0042$. At that metallicity, the BBH would not merge within the age of the Universe unless dynamically assisted by a tertiary (Sect. 3.3), and its merger timescale can be of the order of the ZLK timescale (Fig. 8).
The effects of a tertiary companion around a BBH might be varied and are uncertain. Dorozsmai et al. (2024) demonstrated that triples with CHE binaries can experience tertiary mass transfer onto a BBH, which according to Moreno Méndez et al. (2023) can result in mass growth via hypercritical accretion. Subsequently, Kummer et al. (2025) examined these configurations in greater detail, discovering that gas accretion onto the inner BBH can facilitate drag-assisted mergers, potentially resulting in an electromagnetic counterpart (see also Kıroğlu et al., 2025, for an EM counterpart to a BBH collision with a massive star). In addition, here we find that some BBH mergers can promptly occur purely through dynamical assistance via the ZLK mechanism. In close triples that lead to prompt BBH mergers, the dynamics may fall within a non-adiabatic regime (e.g. Antonini et al., 2018; Rodriguez & Antonini, 2018; Fragione & Kocsis, 2020). In this regime, the black hole spins evolve too slowly to follow the rapid ZLK oscillations through relativistic spin-orbit coupling. When a BBH merges, it losses some mass-energy through gravitational waves, and depending on the spin configuration, it may receive a recoil kick. BBHs formed from CHE evolution are believed to have notably high spins, unlike the typically non-rotating black holes from other evolutionary channels in isolated binary evolution (Marchant et al., 2024). For isolated CHE binaries, these spins are thought to be aligned; however, it is not clear how black hole formation and precession might affect the direction of these spins (e.g. Tauris, 2022). Depending on the alignment of these spins, the recoil kick for the black hole remnant could range from $\sim 10-1000$ km/s. Such a recoil kick could propel the black hole through surrounding material lost via winds or during black hole formation (e.g. Burrows et al., 2023) during the late stages of evolution, as well as prompt interactions with the tertiary if it is a stellar companion. Therefore, we highlight these ZLK-assisted BBH mergers as potential candidates for joint gravitational-wave and electromagnetic detection.
@@ 4.2.1 Stellar evolution
Our analysis examines two stellar models with the same mass, but different metallicities. The evolution of massive stars is influenced by various uncertainties, including convection, mixing, mass loss, and magnetic fields. Here, we emphasise two main caveats: stellar expansion and collapse.
@@
In this paper, we focus on the CHE scenario, where rotational mixing allows the star to remain compact throughout the main sequence and leads to contraction afterwards. Vigna-Gómez et al. (2021) noted that any compact star can play a significant role in the evolution of tight triples, including stars in metal-free (Population III) environments (e.g. Marigo et al., 2001) or within specific mass–metallicity regimes (e.g. Shenar et al., 2020). We adhere this idea and propose that prompt (stellar and BBH) mergers are likely to occur in systems where the component stars of the inner binary do not significantly expand after the main sequence, regardless of the physical mechanism behind this.
With respect to stellar collapse, we assume that our $\approx 43.8\ M_{\odot}$ carbon-oxygen core will form a black hole (e.g. Heger et al., 2003; Fryer et al., 2012). We further assume that this black hole is created through direct collapse, without any baryonic mass ejection or associated explosion. Energy is lost exclusively via neutrinos, resulting in a mass decrement of $\sim 0.1\ M_{\odot}$ and no natal kick, consistent with some observed systems (e.g. Vigna-Gómez et al., 2024, and references therein), including triples (e.g. Burdge et al., 2024; Shariat et al., 2025a). However, it is important to acknowledge the possibility of mass ejection during black hole formation, as well as a natal kick. Mass ejection could contribute to the energy reservoir relevant for an electromagnetic counterpart to the BBH merger (e.g. de Mink & King, 2017; Burrows et al., 2023). A natal kick could affect the orbits and their dynamics. These effects are not considered in this study.
@@ 4.2.2 Binary evolution
In our analysis, the role of tides in the inner binary is crucial (Sect. 2.3). Tides influence the evolution of binary and triple systems through various effects, such as rotation, tidal distortion, and tidal dissipation (Sect. 2.3.4).
@@
Besides rotation, the presence of a close companion can further tidally distort the star, enhancing mixing efficiency compared to models that consider only rotation (Hastings et al., 2020). Additionally, in a tidally deformed short-period binary a non-dissipative tidal bulge may form. This bulge can lead to apsidal precession, which suppresses the ZLK mechanism (Sect. 2.2.2). Our default CHE binary is not only a rotating, short-period binary but also a contact system. In contact binaries, tidal distortion can cause a radius discrepancy of up to 5% (Fabry et al., 2022); however, this discrepancy does not significantly impact our calculations.
In our framework for estimating the maximum eccentricity (Sect. 2.2.2), we assume that tides influence the system through the potential created by the non-dissipative tidal bulge, with no energy dissipation occurring in the system. However, in binaries with orbital periods less than approximately 10 days, tides serve as an efficient mechanism for energy dissipation. In this paper, we focused on the dynamical tide, which is suitable for stars with a radiative envelope, as the primary tidal dissipation mechanism (Sect. 2.3); we estimated $\tau_{\rm{circ}}\approx 1500\ \rm yr$ and $\tau_{\rm{dyn}}\approx 23\ \rm yr$. Within the formalism for the dynamical tide, there are major uncertainties in the value of the second-order tidal coefficient $E_{2}$ (e.g. Siess et al., 2013; Kushnir et al., 2017), which is inversely proportional to $\tau_{\rm{circ}}$ (Eq. (12)). Different models for of $E_{2}$ can lead to shorter ($\tau_{\rm{circ}}\approx 185$ yr and $\tau_{\rm{dyn}}\approx 3$ yr, Qin et al. 2018) or longer ($\tau_{\rm{circ}}\approx 2400$ yr and $\tau_{\rm{dyn}}\approx 37$ yr, Yoon et al. 2010) estimates for the tidal timescale at the ZAMS and throughout the stellar lifetime. These different parametrisations can either constrain or expand the parameter space where the ZLK mechanism influences the eccentricity of the inner binary.
Recently, Sciarini et al. (2024) addressed how the formulae from Hurley et al. (2002), which are widely used to estimate the synchronisation timescales via the dynamical tide, are inconsistent with the original model from Zahn (1977). They propose an alternative formula to compute the circularisation timescale via the dynamical tide. For nearly synchronous, equal-mass binaries, such as those in our Fiducial model, the synchronisation timescale should be considerably longer than the estimate provided in Sect. 2.3. This implies that tidal dissipation in the inner binary becomes negligible (Sect. 2.3.4), unless the binary spins become significantly asynchronous due to perturbations by the tertiary.
Finally, we briefly discuss the role of wind mass loss, which becomes increasingly important at high metallicities (Fig. 9). For the component stars, mass loss through winds ultimately influences the remnant mass of the black hole. More crucially, this mass loss alters the orbital separation — a key factor in determining whether a BBH can merge within the age of the Universe (Peters, 1964). Our model assumes that each star loses mass via a fast isotropic wind that carries away the specific angular momentum of the mass-losing star. This mechanism, often termed the ‘Jeans mode’ or ‘fast winds’ mode (Huang, 1963), leads to orbital widening proportional to the amount of mass lost. Consequently, wide ($\sim 1000$ d) triples at moderately-low metallicity (Fig. 8) do not result in BBH mergers, unlike similar initial configurations at low metallicity (Fig. 7). While fast winds are appropriate for massive, close binaries, slow winds could induce drag and limit the orbital widening (e.g. Schrøder et al., 2021). Additionally, the orbital evolution of the inner binary might modify the triple dynamics, potentially triggering a dynamical instability or affecting ZLK cycles.
@@ 4.2.3 Triple evolution
In the study of multiple star systems, it is often assumed that the component stars are coeval. This assumption was applied in our analysis. However, simulations indicate that some triples form after the inner binary has assembled, either as part of the same starburst (e.g. Bonnell & Bate, 1994) or through a dynamical capture. In our tight triple systems, the ZLK timescale ($\tau_{\rm ZLK}\lesssim 10^{4}$ yr) is very short with respect to the stellar lifetime ($\sim 10^{6}$ yr). Consequently, for a coeval triple system, a merger near the ZAMS would effectively correspond to a merger in the pre-main-sequence phase. Alternatively, the merger might occur at a much later stage, after the triple system has formed and achieved a dynamical steady-state equilibrium. Only then would it evolve in a manner more akin to the scenario considered in our formalism. In such cases, the merger would not occur around the ZAMS but shortly after the triple assembled.
Once the triple system has formed, there are additional considerations to keep in mind. In our analysis, we have assumed that the mass ratio is close to one and that the outer orbit is circular. This assumption for the mass ratio is common in the literature on CHE binaries (Marchant et al., 2016; Mandel & de Mink, 2016). However, massive contact binaries are often observed with mass ratios that differ from unity (see the sample presented in Menon et al., 2021). While Fabry et al. (2023) determined that energy transfer in massive contact binaries alters the evolution of the mass ratio and can prolong the lifespan of a contact system with unequal masses, Fabry et al. (2025) showed that this effect is small at a population level.
The assumption about a circular outer orbit might not be representative for hierarchical triple systems (e.g. Hwang et al., 2022). The two tight high-mass triples, TIC 470710327 (Eisner et al., 2022) and TIC 290061484 (Kostov et al., 2024), exhibit outer eccentricities of $e_{\rm out}=0.3$ and $e_{\rm out}=0.2$, respectively. Also, the tertiary around a massive binary may have an eccentric orbit due to the apsidal precession resonance (e.g. Liu & Lai, 2024). For systems with non-equal mass binaries and non-zero outer eccentricity, the octupole effect may become important (Naoz, 2016). This tends to widen the inclination window for large eccentricity excitation. However, the analytic expression for $e_{\rm max}$ given by Eq. (6) remains valid. Thus, because of the effect of short-range forces due to tides, the maximum eccentricity cannot exceed $e_{\rm max}$ even when the octupole potential is significant (Liu et al., 2015).
@@ 4.3 Constraining tidal theory
There are significant uncertainties in our understanding of tidal theory, particularly in massive, short-period binary systems. These uncertainties primarily stem from the standard model of the dynamical tide (Zahn, 1975, 2008), which describes how stellar oscillation modes are excited by a periodic tidal potential. A key parameter in this model, the $E_{2}$ coefficient (Sect. 2.3), is highly sensitive to the definition of the core-envelope boundary (e.g. Siess et al., 2013; Kushnir et al., 2017), making it a crucial but uncertain factor in tidal evolution.
@@
The interplay between the inner and outer binaries remains highly uncertain. Short-range forces, including rotation, tides, and GR effects, have been shown to suppress the ZLK mechanism (Liu et al., 2015). In our default CHE binary model, tides are the dominant short-range force responsible for this suppression. However, the structure and evolution of the tidal bulge, which determines the maximum eccentricity of the inner binary and whether it will merge as a stellar or BBH system, are still not well understood.
To refine tidal theory, we propose using future observations of high-mass tight triple systems. The mere detection of such systems could provide empirical constraints on tidal effects. If the orbital configuration suggests that the ZLK mechanism should be active, we can compare cases with and without tidal suppression (ZLK+SA+GR+Tides vs ZLK+SA+GR) to determine whether tides are strong enough to prevent eccentricity growth. If the inner binary remains circular over the ZLK timescale (Sect. 2.3), this would place an upper limit on the tidal dissipation timescale. Additionally, such systems could allow us to solve for $E_{2}$ in specific binary configurations, further improving tidal models. Conversely, if the inner binary exhibits eccentricity evolution, we can compare the observed evolution to our ZLK model to assess the accuracy of our tidal prescriptions, either by determining the value of $\epsilon_{\rm Tide}$ or by quantifying deviations from the expected ZLK+SA+GR solution.
In summary, observations of tight triple systems offer a direct way to test and refine tidal theory. By analysing their orbital configurations and evolutionary pathways, we can place new constraints on tidal dissipation, improve our understanding of how tides suppress eccentricity growth, and ultimately enhance our models of stellar and compact object mergers.
@@ 4.4 Population forecasting for tight triples
Finally, tight triples in which the inner binary forms a BBH present very interesting prospects. These can result in triple systems with a tertiary in an outer orbital period of 100 or 1000 d, detectable with facilities such as Gaia (Gaia Collaboration et al., 2016). Triple configurations that lead to a prompt BBH merger could potentially be detected by the LIGO-Virgo-Kagra detector network, resulting in a binary configuration of an unusually massive black hole orbited by a nearby stellar companion.
@@ 5 Summary and conclusions
We investigated the role of the ZLK mechanism in high-mass triple systems. We focused on ‘tight’ triples, which are composed of a short-period, high-mass stellar binary with a nearby tertiary companion. We explored an illustrative case involving a circular binary with component masses $m_{1}=m_{2}=55\ M_{\odot}$, initial orbital period $P_{\rm in}=1.1$ d at ’low’ ($Z=0.00042$) and ‘moderately low’ ($Z=0.0042$) metallicities (Sect. 2.1 and Appendix A, respectively). This binary was simulated using the detailed stellar evolution software MESA.
@@
The component stars in this binary undergo chemically homogeneous evolution (CHE), resulting in well-mixed stars that remain compact throughout the main sequence and contract afterwards (Fig. 1 and 9). Based on our numerical model, we examined the impact of a tertiary with a mass range between $1\leq m_{3}/M_{\odot}\leq 100$ and an outer orbital period range between $1\leq P_{\rm out}/\rm{d}\leq 1000$, assuming the initial mutual inclination is uniformly distributed in the range of $-1\leq\cos(i_{0})\leq 1$.
@@
We utilised the analytic results from Liu et al. (2015) and Mangipudi et al. (2022) to determine the maximum eccentricity that the inner binary can achieve in different triple configurations (Sect. 2.2 and Fig. 5). We evaluated whether the ZLK mechanism could cause the component stars of the binary to fill their outer Lagrangian points. If this occurs, we consider the inner binary to result in a stellar merger (Sect. 3.2). We find three key evolutionary stages for mergers: near the ZAMS, shortly after the end of the main sequence, or after the formation of a BBH (Figs. 7 and 8). Our concluding remarks are as follows.
Main-sequence mergers.
@@
During the main sequence, strong tidal bulges in close binaries significantly damp the maximum eccentricity produced by the ZLK mechanism (Figs. 3 and 5). Nonetheless, mergers can occur even at moderately-low eccentricities ($e\gtrsim 0.3$), with an occurrence fraction of $f_{\rm{merger}}\approx 0.77$ for optimal inclinations. These mergers are possible in systems with outer periods of up to $\approx 70$ d at low metallicity (Fig. 6) and up to $\approx 100$ d at moderately-low metallicity (Fig. 10). Given the ZLK timescales (Fig. 4), these hydrogen-rich mergers must happen near the ZAMS for coeval triples or shortly after the formation of the triple for non-coeval systems.
Post-main-sequence mergers.
@@
After the main sequence, CHE stars rapidly contract and the effect of tides weakens. At this stage, mergers can occur at high eccentricities ($e\gtrsim 0.8$), and occur less frequently at a fraction of $f_{\rm merger}\approx 0.57$ for optimal inclinations. These mergers are helium-rich and can occur in systems with outer periods of up to $\approx 100$ d at low metallicity (Fig. 6) and up to $\approx 270$ d at moderately-low metallicity (Fig. 10).
Binary black hole mergers.
@@
For triples in which the inner binary does not undergo a stellar merger, a BBH may form. In BBHs, tides are no longer present, and the ZLK mechanism is largely suppressed by GR effects. The merger of BBHs can be facilitated by a tertiary companion with an orbital period of up to a few thousand days, and in favourable orbital configurations the time-to-coalescence may be as short as the ZLK timescale (Fig. 7 and 8). This results in prompt mergers shortly after the formation of a BBH within a tight triple system, which contrasts with the generally slower coalescence expected for BBHs formed in isolation. These BBH mergers occur near recently ejected gas (from stellar winds or following core collapse) and possibly close to another star. This makes them promising candidates for electromagnetic counterparts and probes of environmental effects in gravitational-wave sources.
Orbital configurations leading to mergers.
@@
Contrary to the test particle limit, maximum eccentricity is not always achieved at a $90$ deg mutual inclination. Instead, it depends on the ratio of the inner to outer orbital angular momenta ($\eta:=L_{\rm in}/L_{\rm out}$, Fig. 2). Moreover, when the tertiary is less massive than the inner component stars of the binary, the maximum eccentricity is reached at largely retrograde configurations (Fig. 6 and 10). This behaviour is independent of the nature of the component stars of the inner binary; it should also occur in stars that do not evolve via CHE.
Outlook.
@@
The dynamics of tight, high-mass stellar triples are complex. When the tertiary star is more massive than either component stars of the inner binary, the evolution of the system approximates the test particle limit. However, in this regime, stellar evolution (particularly mass transfer from the tertiary) can play a crucial role. Conversely, if the tertiary is less massive than either inner companions, evolution is dominated by the inner binary, and even significantly inclined systems may experience reduced dynamical effects due to being far from the test particle limit. On a population level, an important question remains as to which tight triples are eventually reduced to binaries and which remain intact as triples, noting that the latter stand as candidate progenitors of merging BBHs.